OpenStudy (anonymous):

using the properties of logarithms, expand the expression: ln (4x^5-x-1)(square root x-7)/(x^2+1)^3

OpenStudy (anonymous):

The rules you want are log(a*b)=log(a)+log(b); log(c/d)=log(c)-log(d); and log(a^n)=n*log(a). It depends how far you want to simplify but I get \[\ln(4x^5-x-1)+(1/2)\ln(x-7)-3\ln(x^2+1)\]

OpenStudy (anonymous):

yeah that was my problem i didnt know how far to go...thanks so much